# Large Language Models for Scientific Idea Generation: A Creativity-Centered Survey

### v1

@@ 8.1 Embracing the Open-Ended Nature of Scientific Discovery
Concepts from open-ended learning offer useful inspiration for future directions. Novelty search demonstrated that progress can emerge by prioritizing behavioral novelty over explicit objectives (lehman2011novelty). Quality-diversity methods such as MAP-Elites show how exploration can be structured to discover diverse, high-performing solutions across a wide range of niches (mouret2015mapelites). Co-evolutionary frameworks such as POET (wang2019poet) and hard-exploration strategies like Go-Explore (ecoffet2019goexplore; ecoffet2021firstreturnthenexplore) illustrate how agents can sustain long-term innovation by generating new tasks alongside new solutions, or by systematically revisiting and extending prior stepping stones. At larger scales, environments like XLand demonstrate that procedurally generated, open-ended multi-task worlds can produce generally capable agents through continual adaptation (team2021openended). Earlier theoretical work such as POWERPLAY (schmidhuber2011powerplay) also emphasized continual self-invented tasks as a foundation for unbounded discovery.
@@
8.2 Establishing Standardized Benchmarks and Transparent Evaluation Frameworks
@@
Evaluation remains the most critical bottleneck for LLM-based scientific ideation. At present, nearly every study introduces its own pipeline for assessing novelty, diversity, or feasibility, which makes systematic comparison across approaches nearly impossible (ismayilzada2025creativityaiprogresseschallenges; gupta2024scind; ruan2025liveideabenchevaluatingllmsdivergent). In terms of creativity theory, this corresponds to the Product dimension: creative outcomes are the dependent variables by which novelty, usefulness, and correctness are judged (kozbelt2010theories). Yet current methods lack standardized and transparent metrics for these criteria, leaving evaluations subjective and inconsistent. Automated metrics such as embedding-based novelty scores (wang2025enablingaiscientistsrecognize), diversity indices (friedman2023vendiscorediversityevaluation), or literature-grounded overlap detection (shahid2025literaturegroundednoveltyassessmentscientific) are valuable but insufficient on their own. More importantly, human evaluation practices remain opaque: many papers fail to clearly report annotator expertise, evaluation protocols, or inter-rater reliability, leaving results subjective and difficult to reproduce (gehrmann2022repairingcrackedfoundationsurvey; laskar2024systematicsurveycriticalreview).
@@
A crucial step forward is the development of shared benchmarks and transparent evaluation pipelines. These should span multiple domains (e.g., biology, chemistry, mathematics, social sciences) and multiple levels of abstraction, from hypothesis plausibility to experimental feasibility. Human evaluation in particular needs greater rigor: standardized reporting of evaluator backgrounds, annotation guidelines, and agreement metrics should become community norms. Beyond laboratory-style annotation, peer-review–style evaluations hold particular promise. For instance, the AI Scientist project submitted an AI-generated paper to a workshop as a form of external validation (lu2024aiscientistfullyautomated), demonstrating how conference-style review could provide structured and credible human feedback. Similar efforts—whether via workshops, structured review panels, or crowd-sourced expert collectives—could provide scalable, transparent, and standardized human-in-the-loop evaluation.
@@
By establishing community-accepted benchmarks and rigorous human evaluation frameworks, the field can move beyond isolated proof-of-concept systems toward results that are comparable, reproducible, and trusted by the broader scientific community.
@@ 8.3 Building Domain-Rich Simulators for Feasible and Scalable Training
The role of simulators in accelerating progress is well established in other areas of AI. In robotics and control, physics-based simulators such as MuJoCo (6386109) or Isaac Gym (makoviychuk2021isaac) have enabled breakthroughs in RL by allowing agents to practice millions of interactions before real-world deployment. In scientific discovery, similar trends are emerging: AtomAgents integrates alloy design with materials simulators for iterative hypothesis refinement (ghafarollahi2024atomagentsalloydesigndiscovery), while Robin uses in silico biological assays and robotic wet-labs to close the loop between hypothesis generation and experimental validation (ghareeb2025robinmultiagentautomatingscientific). More abstractly, multi-agent simulation environments such as Casevo (jiang2024casevocognitiveagentssocial) show how agent societies can serve as dynamic testbeds for evaluating ideation in complex social or political domains.
@@ 8.4 Beyond Next-Token Prediction: Architectural Bottlenecks for Creativity
A final challenge concerns the architectural limits of current LLMs, which are fundamentally trained on the next-token prediction objective. Although this paradigm produces fluent language and strong surface-level reasoning, it inherently biases models toward local coherence and incremental extrapolation, discouraging bold conceptual leaps that scientific discovery often requires (bender2021dangers; marcus2022deep). This tendency toward “safe” continuations can stifle creativity and prevent models from venturing into genuinely novel scientific hypotheses. From the perspective of the 4Ps framework, this reflects a limitation of the Person dimension: the intrinsic capacities of the idea-generating system are constrained by its training objectives and inductive biases. Emerging research points to several promising alternatives Multi-token prediction frameworks, such as Roll the Dice and Look Before You Leap (nagarajan2025rolldicelook), seek to reduce the local myopia of auto-regressive decoding by predicting multiple tokens jointly, improving global coherence. State-space sequence models such as Mamba demonstrate long-context stability and efficient memory handling, offering potential for sustained scientific reasoning (gu2024mambalineartimesequencemodeling; halloran2024mambastatespacemodelslyapunovstable). Meanwhile, Diffusion-based language models (nie2025largelanguagediffusionmodels; li2022diffusionlm) depart entirely from the auto-regressive paradigm by framing text generation as iterative denoising, which can encourage more flexible and globally optimized exploration. Hybrid approaches that combine these paradigms with transformers may provide the best path forward, balancing local fluency with global creativity.

### v2

@@ 9.1 Embracing the Open-Ended Nature of Scientific Discovery
Concepts from open-ended learning offer useful inspiration for future directions. Novelty search demonstrated that progress can emerge by prioritizing behavioral novelty over explicit objectives (Lehman and Stanley, 2011). Quality-diversity methods such as MAP-Elites show how exploration can be structured to discover diverse, high-performing solutions across a wide range of niches (Mouret and Clune, 2015). Co-evolutionary frameworks such as POET (Wang et al., 2019) and hard-exploration strategies like Go-Explore (Ecoffet et al., 2019; 2021) illustrate how agents can sustain long-term innovation by generating new tasks alongside new solutions, or by systematically revisiting and extending prior stepping stones. At larger scales, environments like XLand demonstrate that procedurally generated, open-ended multi-task worlds can produce generally capable agents through continual adaptation (Team et al., 2021). Earlier theoretical work such as POWERPLAY (Schmidhuber, 2011) also emphasized continual self-invented tasks as a foundation for unbounded discovery.
@@
9.2 Establishing Standardized Benchmarks and Robust Evaluation Frameworks
@@
Evaluation remains the most critical bottleneck for LLM-based scientific ideation. At present, nearly every study introduces its own pipeline for assessing novelty, diversity, or feasibility, which makes systematic comparison across approaches nearly impossible (Ismayilzada et al., 2025b; Gupta et al., 2024; Ruan et al., 2024). In terms of creativity theory, this corresponds to the Product dimension: creative outcomes are the dependent variables by which novelty, usefulness, and correctness are judged (Kozbelt et al., 2010). Yet current methods lack standardized and transparent metrics for these criteria, leaving evaluations subjective and inconsistent. Automated metrics such as embedding-based novelty scores (Wang et al., 2025e), diversity indices (Friedman and Dieng, 2023), or literature-grounded overlap detection (Shahid et al., 2025) are valuable but insufficient on their own. More importantly, human evaluation practices remain opaque: many papers fail to clearly report annotator expertise, evaluation protocols, or inter-rater reliability, leaving results subjective and difficult to reproduce (Gehrmann et al., 2022; Laskar et al., 2024).
@@
A crucial step forward is the development of shared benchmarks and transparent evaluation pipelines. These should span multiple domains (e.g., biology, chemistry, mathematics, social sciences) and multiple levels of abstraction, from hypothesis plausibility to experimental feasibility. Human evaluation in particular needs greater rigor: standardized reporting of evaluator backgrounds, annotation guidelines, and agreement metrics should become community norms. Beyond laboratory-style annotation, peer-review–style evaluations hold particular promise. For instance, the AI Scientist project submitted an AI-generated paper to a workshop as a form of external validation (Lu et al., 2024a), demonstrating how conference-style review could provide structured and credible human feedback. Similar efforts—whether via workshops, structured review panels, or crowd-sourced expert collectives—could provide scalable, transparent, and standardized human-in-the-loop evaluation.
Beyond human assessment, future work must also address limitations of LLM-based automated evaluators. Supervised fine-tuned judges can suffer from memorization and overfitting, limiting their ability to generalize to genuinely novel ideas. Recent directions suggest augmenting evaluators with online retrieval and test-time reasoning, allowing judgments to be grounded in up-to-date literature and refined through iterative evaluation rather than static pattern matching(Zhu et al., 2025b). More broadly, no single evaluator is sufficient for all criteria: novelty assessment benefits from retrieval against current research, feasibility is best evaluated via domain-specific executors or simulators, and overall correctness or coherence may be more reliably judged by fine-tuned expert models. This motivates a modular, criterion-specific evaluation paradigm rather than monolithic scoring.
Finally, an important open challenge is spurious learning in trained evaluators, where models may rely on superficial cues (e.g., topic popularity or stylistic novelty) instead of core scientific qualities. Addressing this requires techniques inspired by the spurious correlation mitigation literature, such as balanced or counterfactual sampling, re-weighting of evaluation data, and explicit control over correlated attributes during evaluator training(Yang et al., 2023). Incorporating these strategies is essential for building reliable and fair automated reviewers for scientific ideation.
@@
By establishing community-accepted benchmarks, rigorous human evaluation practices, and robust automated evaluators, the field can move beyond isolated proof-of-concept systems toward results that are comparable, reproducible, and trusted by the broader scientific community.
@@ 9.3 Building Domain-Rich Simulators for Feasible and Scalable Training
The role of simulators in accelerating progress is well established in other areas of AI. In robotics and control, physics-based simulators such as MuJoCo (Todorov et al., 2012) or Isaac Gym (Makoviychuk et al., 2021) have enabled breakthroughs in RL by allowing agents to practice millions of interactions before real-world deployment. In scientific discovery, similar trends are emerging: AtomAgents integrates alloy design with materials simulators for iterative hypothesis refinement (Ghafarollahi and Buehler, 2024), while Robin uses in silico biological assays and robotic wet-labs to close the loop between hypothesis generation and experimental validation (Ghareeb et al., 2025). More abstractly, multi-agent simulation environments such as Casevo (Jiang et al., 2024) show how agent societies can serve as dynamic testbeds for evaluating ideation in complex social or political domains.
@@ 9.4 Beyond Next-Token Prediction: Architectural Bottlenecks for Creativity
A final challenge concerns the architectural limits of current LLMs, which are fundamentally trained on the next-token prediction objective. Although this paradigm produces fluent language and strong surface-level reasoning, it inherently biases models toward local coherence and incremental extrapolation, discouraging bold conceptual leaps that scientific discovery often requires (Franceschelli and Musolesi, 2024; Bender et al., 2021; Marcus, 2022). This tendency toward “safe” continuations can stifle creativity and prevent models from venturing into genuinely novel scientific hypotheses. From the perspective of the 4Ps framework, this reflects a limitation of the Person dimension: the intrinsic capacities of the idea-generating system are constrained by its training objectives and inductive biases. Emerging research points to several promising alternatives Multi-token prediction frameworks, such as Roll the Dice and Look Before You Leap (Nagarajan et al., 2025), seek to reduce the local myopia of auto-regressive decoding by predicting multiple tokens jointly, improving global coherence. State-space sequence models such as Mamba demonstrate long-context stability and efficient memory handling, offering potential for sustained scientific reasoning (Gu and Dao, 2024; Halloran et al., 2024). Meanwhile, Diffusion-based language models (Nie et al., 2025; Li and Liang, 2022) depart entirely from the auto-regressive paradigm by framing text generation as iterative denoising, which can encourage more flexible and globally optimized exploration. Hybrid approaches that combine these paradigms with transformers may provide the best path forward, balancing local fluency with global creativity.